# Video Latent Flow Matching: Optimal Polynomial Projections for Video Interpolation and Extrapolation

### v1

@@
Yang Cao ycao4@wyomingseminary.org. Wyoming Seminary.    Zhao Song magic.linuxkde@gmail.com. Simons Institute for the Theory of Computing, University of California, Berkeley.    Chiwun Yang christiannyang37@gmail.com Sun Yat-sen University.
@@ 1 Introduction
  • We validate our approach by conducting extensive experiments. Our model leverages DiT-XL-2 and is trained on a diverse collection of seven large-scale video datasets, including OpenVid-1M, MiraData, and videos from Pixabay. The results demonstrate strong performance in text-to-video generation, interpolation, and extrapolation, achieving robust and reliable outputs with significant potential for real-world video applications.
@@ 2 Related Work
Theory in Transformer-Based Models. Transformers have become a cornerstone in AI and are widely used in different areas, especially in NLP (Natural Language Process) and CV (Computer Vision). However, understanding the Transformers from a theoretical perspective remains an ongoing challenge. Several works have explored the theoretical foundations and computational complexities of the Transformers [97, 112, 16, 3, 93, 18, 43, 76, 96, 8, 24, 100, 92, 5, 27, 56, 72, 32], focusing on areas such as efficient Transformers [42, 89, 95, 63, 66, 71, 70, 68, 67, 21, 65, 48, 47, 39, 104, 49, 4, 35], optimization [26], and the analysis of emergent abilities [12, 107, 7, 51, 109, 64, 108, 19, 69, 46, 106]. Notably, [112] introduced an algorithm with provable guarantees for approximation of Transformers, [59] proved a lower bound for Transformers based on the Strong Exponential Time Hypothesis, and [3] provided both an algorithm and hardness results for static Transformers computation.
@@ 5.1 Diffusion Transformer (DiT)
Diffusion Transformer [80] is a framework that utilizes Transformers [103] as the backbone for Diffusion Models [41, 88]. Specifically, a Transformer block consists of a multi-head self-attention layer and a feed-forward layer, with both layers having a skip connection. We use ${\sf TF}^{h,m,r}:\mathbb{R}^{n\times d_{0}}\rightarrow\mathbb{R}^{n\times d_{0}}$ to denote a Transformer block. Here $h$ and $m$ are the number of heads and head size in self-attention layer, and $r$ is the hidden dimension in feed-forward layer. Let $X\in\mathbb{R}^{n\times d_{0}}$ be the model input. Then, we have the model output
$~{}{\sf Attn}(X)$
$~{}:=\sum_{i=1}^{h}{\sf Softmax}(XW_{Q}^{i}{W_{K}^{i}}^{\top}X^{\top})\cdot XW_{V}^{i}{W_{O}^{i}}^{\top}+X.$
where $W_{K}^{i},W_{Q}^{i},W_{V}^{i},W_{O}^{i}\in\mathbb{R}^{d_{0}\times m}$. Moreover,
@@
where $W_{1},W_{2}\in\mathbb{R}^{d_{0}\times r},b_{1}\in\mathbb{R}^{r},b_{2}\in\mathbb{R}^{d_{0}}$.
We use Transformer networks with positional encoding $E\in\mathbb{R}^{n\times d_{0}}$ in our work. The transformer networks are then defined as the composition of Transformer blocks
@@ 5.2 Approximation via DiT
Theorem 5.1 (Informal version of Theorem D.7).
@@
Please refer to the proof of Theorem D.7 for the detailed analysis. ∎
@@ 5.3 Interpolation and Extrapolation
Lemma 5.2 (Informal version of Lemma E.3).
@@
Proof sketch of Lemma 5.2 This proof follows from its formal version in Lemma E.3
@@
Theorem 5.3 (Informal version of Theorem E.4).
@@
Please refer to Theorem E.4 for complete proofs. ∎
@@ 5.4 Timescale Robustness
Following [30], we demonstrate that projection onto from latent patches $u_{t}$ is robust to timescales. Formally, the HiPPO-LegS operator is timescale-equivariant: dilating the input $u$ does not change the approximation coefficients $H_{N}$. At the same time, this property is working in the case of the discretized form $\widetilde{u}$. We emphasize that it is crucial to use flow matching to model the latent patches, where whatever the sampling method and frame rate are, it will not greatly harm VLFM’s performance. We give its formal statement below.
Lemma 5.4 (Proposition 3 of [30], informal version of Lemma E.2).
@@ 6.2 Text-to-Video Generation with Arbitrary Frame Rate
In this section, we recover several videos with different frame rates using VLFM with given video captions in the training dataset. We extract $T=0.5$ for demonstrations as Figure 2. In detail, we choose three frame rates for generation $\{8,12,16\}$. As shown, our VLFM performs fairly on text-to-video generation while it requires very small resource that is equivalent to training a new flow matching text-to-image video, which ensures its efficiency. Moreover, we give more results that are generated by VLFM in Appendix A.
@@ 7 Conclusion
Limitations. Since the motivation of this paper focuses on simply and efficiently solving the main goal, it lacks enough exploring each design and how it affects the empirical performance, providing little insights for the follow-ups. Hence, we leave these comprehensive explorations, the study of the most efficient design of VLFM, and its more concise theoretical working mechanism behind as future works.
@@
In the appendix, we first present more experimental text-to-video generation results in Section A. Then we introduce the preliminary in Section B. Next, we illustrate Video Latent Flow Matching formally in Section C. In Section D, we demonstrate the Diffusion Transformer, and finally, in Section E, we present the interpolation and extrapolation of VLFM.
@@ Appendix A More Text-to-Video Generation Results
Appendix B Preliminary
In the preliminary section, we first introduce our notation in the appendix in Section B.1. Then, in Section B.2, we formally define the video-caption data and visual decoder. In Section B.3, we define the latent patches. Section B.4 makes some assumptions which we will use later. Finally, in Section B.5, we list some basic useful facts.
B.1 Notations
@@
B.2 Video-Caption Data
Definition B.1 (Video-caption data pairs and their distribution).
@@ B.2 Video-Caption Data
Definition B.2.
Given a video caption distribution ${\cal V}_{c}$ as Definition B.1. We denote $\Delta t$ as the minimal time unit of measurement in the real world (Planck time). For any $(V,c)\sim{\cal V}_{c}$, we define the discretized form of $V:[0,T]\rightarrow\mathbb{R}^{D}$, which is $\widetilde{V}\in\mathbb{R}^{\frac{T}{\Delta t}\times D}$, and its $\tau$-th row $\forall\tau\in[\frac{T}{\Delta t}]$ is given by:
@@
Definition B.3 (Obtained data in real-world cases).
@@
  • Given a video caption distribution ${\cal V}_{c}$ as Definition B.1.
  • For any $(V,c)\sim{\cal V}_{c}$, we define the discretized form of video $\widetilde{V}$ as Definition B.2.
@@
Definition B.4 (Bijective Visual Decoder).
@@
B.3 Latent Patches Data
Definition B.5.
@@ B.3 Latent Patches Data
  • Given a video caption distribution ${\cal V}_{c}$ as Definition B.1.
  • For any $(V,c)\sim{\cal V}_{c}$, we define the discretized form of video $\widetilde{V}$ as Definition B.2.
  • Let the observation matrix $\Phi:\{0,1\}^{N\times\frac{T}{\Delta t}}$ be defined as Definition B.3.
  • Let the visual decoder function $D:\mathbb{R}^{d}\rightarrow\mathbb{R}^{D}$ be defined as Definition B.4.
@@
  • Given a video caption distribution ${\cal V}_{c}$ as Definition B.1.
  • For any $(V,c)\sim{\cal V}_{c}$, we define the discretized form of video as Definition B.2.
  • Let the observation matrix $\Phi:\{0,1\}^{N\times\frac{T}{\Delta t}}$ be defined as Definition B.3.
  • Let the visual decoder function $D:\mathbb{R}^{d}\rightarrow\mathbb{R}^{D}$ be defined as Definition B.4.
@@ B.4 Assumptions
  • Given a video caption distribution ${\cal V}_{c}$ as Definition B.1.
  • For any $(V,c)\sim{\cal V}_{c}$, we define the discretized form of video as Definition B.2.
  • Let the observation matrix $\Phi:\{0,1\}^{N\times\frac{T}{\Delta t}}$ be defined as Definition B.3.
  • Let the visual decoder function $D:\mathbb{R}^{d}\rightarrow\mathbb{R}^{D}$ be defined as Definition B.4.
  • Let the ideal version of the sequence of latent patches $u\in\mathbb{R}^{\frac{T}{\Delta t}\times d}$ be defined as Definition B.5.
@@
$\lim_{\Delta t\rightarrow 0}\frac{u_{\tau+1}^{(i-1)}-u_{\tau}^{(i-1)}}{\Delta t},\forall i\in[k],\tau\in[\frac{T}{\Delta t}].$
Assumption B.8.
@@
  • Let the visual decoder function $D:\mathbb{R}^{d}\rightarrow\mathbb{R}^{D}$ be defined as Definition B.4.
@@
  • Given a video caption distribution ${\cal V}_{c}$ as Definition B.1.
  • For any $(V,c)\sim{\cal V}_{c}$, we define the discretized form of video as Definition B.2.
  • Let the observation matrix $\Phi:\{0,1\}^{N\times\frac{T}{\Delta t}}$ be defined as Definition B.3.
  • Let the visual decoder function $D:\mathbb{R}^{d}\rightarrow\mathbb{R}^{D}$ be defined as Definition B.4.
  • Let the ideal version of the sequence of latent patches $u\in\mathbb{R}^{\frac{T}{\Delta t}\times d}$ be defined as Definition B.5.
@@
  • Given a video caption distribution ${\cal V}_{c}$ as Definition B.1.
@@
B.5 Basic Facts
Fact B.11.
@@ B.5 Basic Facts
Appendix C Video Latent Flow Matching
This section, we first introduce the HiPPO Framework and LegS in Section C.1. Then, we formally define the video latent flow in Section C.2. Last, we introduce the training objective of VLFM in Section C.3.
C.1 HiPPO Framework and LegS
Definition C.1.
@@ C.1 HiPPO Framework and LegS
  • Let matrix $A\in\mathbb{R}^{s\times s}$ be defined as Definition C.1.
  • Let matrix $B\in\mathbb{R}^{s\times 1}$ be defined as Definition C.2.
We initialize a matrix $H_{0}\in\mathbb{R}^{d\times s}$ where each of its entries is sampled from Gaussian. Then we define:
@@
C.2 Video Latent Flow
Definition C.5.
@@ C.2 Video Latent Flow
  • Given a video caption distribution ${\cal V}_{c}$ as Definition B.1.
  • For any $(V,c)\sim{\cal V}_{c}$, we define the discretized form of video as Definition B.2.
  • Let the observation matrix $\Phi:\{0,1\}^{N\times\frac{T}{\Delta t}}$ be defined as Definition B.3.
  • Let the visual decoder function $D:\mathbb{R}^{d}\rightarrow\mathbb{R}^{D}$ be defined as Definition B.4.
  • Let the ideal version of the sequence of latent patches $u\in\mathbb{R}^{\frac{T}{\Delta t}\times d}$ be defined as Definition B.5.
  • Let the real-world version of the sequence of latent patches $\widetilde{u}\in\mathbb{R}^{N\times d}$ be defined as Definition B.6.
  • Let $H_{N}\in\mathbb{R}^{d\times s}$ be defined as Definition C.3.
  • Let the function of polynomials $g(t)$ be defined as Definition C.4.
@@
  • Given a video caption distribution ${\cal V}_{c}$ as Definition B.1.
  • For any $(V,c)\sim{\cal V}_{c}$, we define the discretized form of video as Definition B.2.
  • Let the observation matrix $\Phi:\{0,1\}^{N\times\frac{T}{\Delta t}}$ be defined as Definition B.3.
  • Let the visual decoder function $D:\mathbb{R}^{d}\rightarrow\mathbb{R}^{D}$ be defined as Definition B.4.
  • Let the ideal version of the sequence of latent patches $u\in\mathbb{R}^{\frac{T}{\Delta t}\times d}$ be defined as Definition B.5.
  • Let the real-world version of the sequence of latent patches $\widetilde{u}\in\mathbb{R}^{N\times d}$ be defined as Definition B.6.
  • Let $H_{N}\in\mathbb{R}^{d\times s}$ be defined as Definition C.3.
  • Let the function of polynomials $g(t)$ be defined as Definition C.4.
@@
  • Given a video caption distribution ${\cal V}_{c}$ as Definition B.1.
  • For any $(V,c)\sim{\cal V}_{c}$, we define the discretized form of video as Definition B.2.
  • Let the observation matrix $\Phi:\{0,1\}^{N\times\frac{T}{\Delta t}}$ be defined as Definition B.3.
  • Let the visual decoder function $D:\mathbb{R}^{d}\rightarrow\mathbb{R}^{D}$ be defined as Definition B.4.
  • Let the ideal version of the sequence of latent patches $u\in\mathbb{R}^{\frac{T}{\Delta t}\times d}$ be defined as Definition B.5.
  • Let the real-world version of the sequence of latent patches $\widetilde{u}\in\mathbb{R}^{N\times d}$ be defined as Definition B.6.
  • Let $H_{N}\in\mathbb{R}^{d\times s}$ be defined as Definition C.3.
  • Let the function of polynomials $g(t)$ be defined as Definition C.4.
  • Let the time-dependent mean of Gaussian distribution $\mu_{t}(\widetilde{u})$ be defined as Definition C.5.
  • Let the time-dependent standard deviation $\sigma_{t}(\widetilde{u})$ be defined as Definition C.6.
@@
  • Given a video caption distribution ${\cal V}_{c}$ as Definition B.1.
  • For any $(V,c)\sim{\cal V}_{c}$, we define the discretized form of video as Definition B.2.
  • Let the observation matrix $\Phi:\{0,1\}^{N\times\frac{T}{\Delta t}}$ be defined as Definition B.3.
  • Let the visual decoder function $D:\mathbb{R}^{d}\rightarrow\mathbb{R}^{D}$ be defined as Definition B.4.
  • Let the ideal version of the sequence of latent patches $u\in\mathbb{R}^{\frac{T}{\Delta t}\times d}$ be defined as Definition B.5.
  • Let the real-world version of the sequence of latent patches $\widetilde{u}\in\mathbb{R}^{N\times d}$ be defined as Definition B.6.
  • Let $H_{N}\in\mathbb{R}^{d\times s}$ be defined as Definition C.3.
  • Let the function of polynomials $g(t)$ be defined as Definition C.4.
  • Let the time-dependent mean of Gaussian distribution $\mu_{t}(\widetilde{u})$ be defined as Definition C.5.
  • Let the time-dependent standard deviation $\sigma_{t}(\widetilde{u})$ be defined as Definition C.6.
@@
C.3 Training Objective
Definition C.9.

### v2

@@
Yang Cao ycao4@wyomingseminary.org. Wyoming Seminary.    Zhao Song magic.linuxkde@gmail.com. Simons Institute for the Theory of Computing, University of California, Berkeley.    Chiwun Yang christiannyang37@gmail.com. Sun Yat-sen University.
@@ 1 Introduction
  • We validate our approach by conducting extensive experiments in Section 6. Our model leverages DiT-XL-2 and is trained on a diverse collection of seven large-scale video datasets, including OpenVid-1M, MiraData, and videos from Pixabay. The results demonstrate strong performance in text-to-video generation, interpolation, and extrapolation, achieving robust and reliable outputs with significant potential for real-world video applications.
@@ 2 Related Work
Theory in Transformer-Based Models. Transformers have become a cornerstone in AI and are widely used in different areas, especially in NLP (Natural Language Process) and CV (Computer Vision). However, understanding the Transformers from a theoretical perspective remains an ongoing challenge. Several works have explored the theoretical foundations and computational complexities of the Transformers [100, 115, 16, 3, 96, 18, 46, 79, 99, 8, 25, 103, 95, 5, 28, 59, 75, 35], focusing on areas such as efficient Transformers [45, 92, 98, 66, 69, 74, 73, 71, 70, 21, 68, 51, 50, 42, 107, 52, 4, 38], optimization [27, 22], and the analysis of emergent abilities [12, 110, 7, 54, 112, 67, 111, 19, 72, 49, 109, 30]. Notably, [115, 29] introduced an algorithm with provable guarantees for approximation of Transformers, [62] proved a lower bound for Transformers based on the Strong Exponential Time Hypothesis, and [3] provided both an algorithm and hardness results for static Transformers computation.
@@ 5.1 Diffusion Transformer (DiT)
Diffusion Transformer [83] is a framework that utilizes Transformers [106] as the backbone for Diffusion Models [44, 91]. Specifically, a Transformer block consists of a multi-head self-attention layer and a feed-forward layer, with both layers having a skip connection. We use ${\sf TF}^{h,m,r}:\mathbb{R}^{n\times d_{0}}\rightarrow\mathbb{R}^{n\times d_{0}}$ to denote a Transformer block. Here $h$ and $m$ are the number of heads and head size in self-attention layer, and $r$ is the hidden dimension in feed-forward layer. Let $X\in\mathbb{R}^{n\times d_{0}}$ be the model input. Then, we have the model output:
${\sf Attn}(X):=\sum_{i=1}^{h}{\sf Softmax}(XW_{Q}^{i}{W_{K}^{i}}^{\top}X^{\top})\cdot XW_{V}^{i}{W_{O}^{i}}^{\top}+X,$
where the projection weights $W_{K}^{i},W_{Q}^{i},W_{V}^{i},W_{O}^{i}\in\mathbb{R}^{d_{0}\times m}$. Moreover,
@@
where the projection weights $W_{1},W_{2}\in\mathbb{R}^{d_{0}\times r}$, bias $b_{1}\in\mathbb{R}^{r},b_{2}\in\mathbb{R}^{d_{0}}$, and $\phi$ is usually considered as the ReLU activated function.
In our work, we use Transformer networks with positional encoding $E\in\mathbb{R}^{n\times d_{0}}$. The transformer networks are then defined as the composition of Transformer blocks:
@@ 5.2 Approximation via DiT
Theorem 5.1 (Informal version of Theorem E.7).
@@
Please refer to the proof of Theorem E.7 for the detailed analysis. ∎
@@ 5.3 Interpolation and Extrapolation
Lemma 5.2 (Informal version of Lemma F.3).
@@
Proof sketch of Lemma 5.2 This proof follows from its formal version in Lemma F.3
@@
Theorem 5.3 (Informal version of Theorem F.4).
@@
Please refer to Theorem F.4 for complete proofs. ∎
@@
  • Information. Besides, a sub-linear factor $\sqrt{\frac{T}{\Delta t}-N}$, which stands for the obtained information about the continuous video, is vital as well for interpolation and extrapolation on data in distribution.
@@ 5.4 Timescale Robustness
Following [33], we demonstrate that projection onto latent patches $u_{t}$ is robust to timescales. Formally, the HiPPO-LegS operator is timescale-equivariant: dilating the input $u$ does not change the approximation coefficients $H_{N}$. At the same time, this property is working in the case of the discretized form $\widetilde{u}$. We emphasize that it is crucial to use flow matching to model the latent patches, where whatever the sampling method and frame rate are, it will not greatly harm VLFM’s performance. We give its formal statement below.
Lemma 5.4 (Proposition 3 of [33], informal version of Lemma F.2).
@@ 6.2 Text-to-Video Generation with Arbitrary Frame Rate
In this section, we recover several videos with different frame rates using VLFM with given video captions in the training dataset. We extract $T=0.5$ for demonstrations as Figure 2. In detail, we choose three frame rates for generation $\{8,12,16\}$. As shown, our VLFM performs fairly on text-to-video generation while it requires very small resource that is equivalent to training a new flow matching text-to-image video, which ensures its efficiency. Moreover, we give more results that are generated by VLFM in Appendix A and B.
@@ 7 Conclusion
Limitations. Since the motivation of this paper focuses on simply and efficiently solving the main goal, it lacks enough exploring each design and how it affects the empirical performance, providing little insights for the follow-ups. Hence, we leave these comprehensive explorations, and its more concise theoretical working mechanism behind as future works. On the other hand, although VLFM simplifies the video modeling process, it necessitates additional computational consumption concerning the combination of the visual decoder part and the flow matching part at the inference stage. We also leave such exploration to a more efficient inference method as a future direction.
@@
In the appendix, we present more experimental text-to-video generation results in Appendix A and more interpolation and extrapolation results in Appendix B. Then we introduce the preliminary in Appendix C. Next, we illustrate Video Latent Flow Matching formally in Appendix D. In Appendix E, we demonstrate the Diffusion Transformer, and finally, in Appendix F, we present the interpolation and extrapolation of VLFM.
@@ Appendix A More Text-to-Video Generation Results
Appendix B More Interpolation and Extrapolation Results
We give more results of interpolation and extrapolation of VLFM in Figure 6.
Figure 6: Interpolation and Extrapolation of VLFM.
Appendix C Preliminary
In the preliminary section, we first introduce our notation in the appendix in Appendix C.1. Then, in Appendix C.2, we formally define the video-caption data and visual decoder. In Appendix C.3, we define the latent patches. Appendix C.4 makes some assumptions which we will use later. Finally, in Appendix C.5, we list some basic useful facts.
C.1 Notations
@@
C.2 Video-Caption Data
Definition C.1 (Video-caption data pairs and their distribution).
@@ C.2 Video-Caption Data
Definition C.2.
Given a video caption distribution ${\cal V}_{c}$ as Definition C.1. We denote $\Delta t$ as the minimal time unit of measurement in the real world (Planck time). For any $(V,c)\sim{\cal V}_{c}$, we define the discretized form of $V:[0,T]\rightarrow\mathbb{R}^{D}$, which is $\widetilde{V}\in\mathbb{R}^{\frac{T}{\Delta t}\times D}$, and its $\tau$-th row $\forall\tau\in[\frac{T}{\Delta t}]$ is given by:
@@
Definition C.3 (Obtained data in real-world cases).
@@
  • Given a video caption distribution ${\cal V}_{c}$ as Definition C.1.
  • For any $(V,c)\sim{\cal V}_{c}$, we define the discretized form of video $\widetilde{V}$ as Definition C.2.
@@
Definition C.4 (Bijective Visual Decoder).
@@
C.3 Latent Patches Data
Definition C.5.
@@ C.3 Latent Patches Data
  • Given a video caption distribution ${\cal V}_{c}$ as Definition C.1.
  • For any $(V,c)\sim{\cal V}_{c}$, we define the discretized form of video $\widetilde{V}$ as Definition C.2.
  • Let the observation matrix $\Phi:\{0,1\}^{N\times\frac{T}{\Delta t}}$ be defined as Definition C.3.
  • Let the visual decoder function $D:\mathbb{R}^{d}\rightarrow\mathbb{R}^{D}$ be defined as Definition C.4.
@@
  • Given a video caption distribution ${\cal V}_{c}$ as Definition C.1.
  • For any $(V,c)\sim{\cal V}_{c}$, we define the discretized form of video as Definition C.2.
  • Let the observation matrix $\Phi:\{0,1\}^{N\times\frac{T}{\Delta t}}$ be defined as Definition C.3.
  • Let the visual decoder function $D:\mathbb{R}^{d}\rightarrow\mathbb{R}^{D}$ be defined as Definition C.4.
@@ C.4 Assumptions
  • Given a video caption distribution ${\cal V}_{c}$ as Definition C.1.
  • For any $(V,c)\sim{\cal V}_{c}$, we define the discretized form of video as Definition C.2.
  • Let the observation matrix $\Phi:\{0,1\}^{N\times\frac{T}{\Delta t}}$ be defined as Definition C.3.
  • Let the visual decoder function $D:\mathbb{R}^{d}\rightarrow\mathbb{R}^{D}$ be defined as Definition C.4.
  • Let the ideal version of the sequence of latent patches $u\in\mathbb{R}^{\frac{T}{\Delta t}\times d}$ be defined as Definition C.5.
@@
$u_{\tau}^{(i)}=\lim_{\Delta t\rightarrow 0}\frac{u_{\tau+1}^{(i-1)}-u_{\tau}^{(i-1)}}{\Delta t},\forall i\in[k],\tau\in[\frac{T}{\Delta t}],$
where, we use $u_{\tau}^{(i)}$ to denote the $i$-th derivation of $u$.
Assumption C.8.
@@
  • Let the visual decoder function $D:\mathbb{R}^{d}\rightarrow\mathbb{R}^{D}$ be defined as Definition C.4.
@@
  • Given a video caption distribution ${\cal V}_{c}$ as Definition C.1.
  • For any $(V,c)\sim{\cal V}_{c}$, we define the discretized form of video as Definition C.2.
  • Let the observation matrix $\Phi:\{0,1\}^{N\times\frac{T}{\Delta t}}$ be defined as Definition C.3.
  • Let the visual decoder function $D:\mathbb{R}^{d}\rightarrow\mathbb{R}^{D}$ be defined as Definition C.4.
  • Let the ideal version of the sequence of latent patches $u\in\mathbb{R}^{\frac{T}{\Delta t}\times d}$ be defined as Definition C.5.
@@
  • Given a video caption distribution ${\cal V}_{c}$ as Definition C.1.
@@
C.5 Basic Facts
Fact C.11.
@@ C.5 Basic Facts
Appendix D Video Latent Flow Matching
This section, we first introduce the HiPPO Framework and LegS in Appendix D.1. Then, we formally define the video latent flow in Appendix D.2. Last, we introduce the training objective of VLFM in Appendix D.3.
D.1 HiPPO Framework and LegS
Definition D.1.
@@ D.1 HiPPO Framework and LegS
  • Let matrix $A\in\mathbb{R}^{s\times s}$ be defined as Definition D.1.
  • Let matrix $B\in\mathbb{R}^{s\times 1}$ be defined as Definition D.2.
We initialize a matrix $H_{0}={\bf 0}_{d\times s}$. Then we define:
@@
D.2 Video Latent Flow
Definition D.5.
@@ D.2 Video Latent Flow
  • Given a video caption distribution ${\cal V}_{c}$ as Definition C.1.
  • For any $(V,c)\sim{\cal V}_{c}$, we define the discretized form of video as Definition C.2.
  • Let the observation matrix $\Phi:\{0,1\}^{N\times\frac{T}{\Delta t}}$ be defined as Definition C.3.
  • Let the visual decoder function $D:\mathbb{R}^{d}\rightarrow\mathbb{R}^{D}$ be defined as Definition C.4.
  • Let the ideal version of the sequence of latent patches $u\in\mathbb{R}^{\frac{T}{\Delta t}\times d}$ be defined as Definition C.5.
  • Let the real-world version of the sequence of latent patches $\widetilde{u}\in\mathbb{R}^{N\times d}$ be defined as Definition C.6.
  • Let $H_{N}\in\mathbb{R}^{d\times s}$ be defined as Definition D.3.
  • Let the function of polynomials $g(t)$ be defined as Definition D.4.
@@
  • Given a video caption distribution ${\cal V}_{c}$ as Definition C.1.
  • For any $(V,c)\sim{\cal V}_{c}$, we define the discretized form of video as Definition C.2.
  • Let the observation matrix $\Phi:\{0,1\}^{N\times\frac{T}{\Delta t}}$ be defined as Definition C.3.
  • Let the visual decoder function $D:\mathbb{R}^{d}\rightarrow\mathbb{R}^{D}$ be defined as Definition C.4.
  • Let the ideal version of the sequence of latent patches $u\in\mathbb{R}^{\frac{T}{\Delta t}\times d}$ be defined as Definition C.5.
  • Let the real-world version of the sequence of latent patches $\widetilde{u}\in\mathbb{R}^{N\times d}$ be defined as Definition C.6.
  • Let $H_{N}\in\mathbb{R}^{d\times s}$ be defined as Definition D.3.
  • Let the function of polynomials $g(t)$ be defined as Definition D.4.
@@
  • Given a video caption distribution ${\cal V}_{c}$ as Definition C.1.
  • For any $(V,c)\sim{\cal V}_{c}$, we define the discretized form of video as Definition C.2.
  • Let the observation matrix $\Phi:\{0,1\}^{N\times\frac{T}{\Delta t}}$ be defined as Definition C.3.
  • Let the visual decoder function $D:\mathbb{R}^{d}\rightarrow\mathbb{R}^{D}$ be defined as Definition C.4.
  • Let the ideal version of the sequence of latent patches $u\in\mathbb{R}^{\frac{T}{\Delta t}\times d}$ be defined as Definition C.5.
  • Let the real-world version of the sequence of latent patches $\widetilde{u}\in\mathbb{R}^{N\times d}$ be defined as Definition C.6.
  • Let $H_{N}\in\mathbb{R}^{d\times s}$ be defined as Definition D.3.
  • Let the function of polynomials $g(t)$ be defined as Definition D.4.
  • Let the time-dependent mean of Gaussian distribution $\mu_{t}(\widetilde{u})$ be defined as Definition D.5.
  • Let the time-dependent standard deviation $\sigma_{t}(\widetilde{u})$ be defined as Definition D.6.
@@
  • Given a video caption distribution ${\cal V}_{c}$ as Definition C.1.
  • For any $(V,c)\sim{\cal V}_{c}$, we define the discretized form of video as Definition C.2.
  • Let the observation matrix $\Phi:\{0,1\}^{N\times\frac{T}{\Delta t}}$ be defined as Definition C.3.
  • Let the visual decoder function $D:\mathbb{R}^{d}\rightarrow\mathbb{R}^{D}$ be defined as Definition C.4.
  • Let the ideal version of the sequence of latent patches $u\in\mathbb{R}^{\frac{T}{\Delta t}\times d}$ be defined as Definition C.5.
  • Let the real-world version of the sequence of latent patches $\widetilde{u}\in\mathbb{R}^{N\times d}$ be defined as Definition C.6.
  • Let $H_{N}\in\mathbb{R}^{d\times s}$ be defined as Definition D.3.
  • Let the function of polynomials $g(t)$ be defined as Definition D.4.
  • Let the time-dependent mean of Gaussian distribution $\mu_{t}(\widetilde{u})$ be defined as Definition D.5.
  • Let the time-dependent standard deviation $\sigma_{t}(\widetilde{u})$ be defined as Definition D.6.
@@
D.3 Training Objective
Definition D.9.